In a certain steam plant the turbine develop 703.72 kJ/kg . The heat supplied to the steam in the boiler is 2800 kJ/kg . The heat rejected by the steam to the cooling water in the condenser is 2100 kJ/kg and the net work of the cycle is -995 kW . Calculate the pump work in kj/kg and steam flow rate.
1
Expert's answer
2020-12-21T11:01:50-0500

Answer

the pump work in kj/kg  is given by

"W=Q_1-Q_2=2800-2100"

"W=700KJ\/kg"

steam flow rate is given as

"\frac{dQ}{dt}=\frac{dW}{dt}=995KJ/kg-sec"